OpenStudy (cookie_2046):

Evaluate 4r(t-v) if r=3, t=5, and v=2. a. 58 b. 9 c. 36 d. 84

OpenStudy (hybrik):

4(3)(5-2) = 12(5-2) = 12(3) = 36, C

OpenStudy (hybrik):

Plug in the points into the values.